9z+8=84.6 can you solve this simple question for me please
Found 2 solutions by vincygrrl22, vidhyak:
Answer by vincygrrl22(7)

9z-8=84.6
9z+84.6+8
9z=92.6
9z/9=92.6/9
z=10.289
*its a round about answer its not exact*

Answer by vidhyak(80)

I saw a mistake in the sign in the answer mentioned, so I'm answering
9z + 8 = 84.6
9z = 84.6 - 8
9z = 76.6
z = 76.6/9
z=8.51